# Subjective-Aligned Dataset and Metric for Text-to-Video Quality Assessment

### v3

@@ 4.2. Video Fidelity Encoder
Type | Models | T2VQA-DB Validation | Sora Testing | ||||||
SROCC | PLCC | KRCC | RMSE | SROCC | PLCC | KRCC | RMSE | ||
zero-shot | CLIPSim (Radford et al., 2021) | 0.1047 | 0.1277 | 0.0702 | 21.683 | 0.2116 | 0.1538 | 0.1406 | 18.316 |
BLIP (Li et al., 2022a) | 0.1659 | 0.1860 | 0.1112 | 18.373 | 0.2126 | 0.1038 | 0.1515 | 18.850 | |
ImageReward (Xu et al., 2024) | 0.1875 | 0.2121 | 0.1266 | 18.243 | 0.0992 | 0.0415 | 0.0748 | 19.494 | |
ViCLIP (Wang et al., 2023b) | 0.1162 | 0.1449 | 0.0781 | 21.655 | 0.2567 | 0.1844 | 0.1734 | 17.982 | |
UMTScore (Liu et al., 2024) | 0.0676 | 0.0721 | 0.0453 | 22.559 | 0.2594 | 0.0840 | 0.1680 | 19.057 | |
finetuned | SimpleVQA (Sun et al., 2022) | 0.6275 | 0.6338 | 0.4466 | 11.163 | 0.0340 | 0.2344 | 0.0237 | 16.687 |
BVQA (Li et al., 2022b) | 0.7390 | 0.7486 | 0.5487 | 15.645 | 0.4235 | 0.2489 | 0.2635 | 17.164 | |
FAST-VQA (Wu et al., 2022) | 0.7173 | 0.7295 | 0.5303 | 10.595 | 0.4301 | 0.2369 | 0.2939 | 17.426 | |
DOVER (Wu et al., 2023e) | 0.7609 | 0.7693 | 0.5704 | 9.8072 | 0.4421 | 0.2689 | 0.2757 | 17.182 | |
Ours | T2VQA | 0.7965 | 0.8066 | 0.6058 | 9.0221 | 0.6485 | 0.3124 | 0.4874 | 16.511 |
@@ 4.4. Quality Regression
LLMs have been proven to be competitive on IQA/VQA tasks (Wu et al., 2024b, 2023f; Zhang et al., 2023c; Wu et al., 2024c). Inspired by them, we also utilize an LLM as the quality regression module in T2VQA. We first design a text instruction prompt, i.e., “Please rate the quality of this video.”, to guide the LLM. The encoded instruction and the fused feature are concatenated as the input of the LLM. Following (Zhang et al., 2023e; Wu et al., 2023f), we supervise the LLM to output one among five ITU-standard (ITU-R, 2002) levels (bad, poor, fair, good, and excellent) to represent the quality of the videos, denoted as <level>. We assign them weights of in order. Since the logit at <level> in LLM is the probability distribution of all tokens, it can be used to represent how the LLM predicts the quality of the video. Therefore, a softmax for each token is calculated and multiplied by its weight. We have the final predicted score as:
@@ 5.1. Implement Details
where is a hyper-parameter for balancing, and is set to during training.
@@
Along with PLCC, we include Spearman’s Rank-Order Correlation Coefficient (SROCC), Kendall’s Rank-order Correlation Coefficient (KRCC), and Root Mean Square Error (RMSE) as performance criteria. SROCC, PLCC, and KRCC indicate the prediction monotonicity, while RMSE measures the prediction accuracy. Better models should have larger SROCC, KRCC, and PLCC scores, but conversely for RMSE. Before calculating PLCC, we follow the same procedure in (Group et al., 2000) to map the objective score to the subject score using a four-parameter logistic function.
@@ 5.2. Performance Comparison
We use CLIPSim (Wu et al., 2021), BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), UMTScore (Liu et al., 2024), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), and DOVER (Wu et al., 2023e) as the reference algorithms. DOVER is considered the SOTA VQA method to date. CLIPSim averages the similarity values between the text and each video frame. We adopt the same operation on BLIP and ImageReward to tune IQA metrics into VQA metrics. ViCLIP and UMTScore are metrics designed for measuring text-video alignment. SimpleVQA, BVQA, FAST-VQA, and DOVER are models designed for general VQA tasks. We use the pre-trained weights of CLIPSim, BLIP, and ImageReward for zero-shot testing, and we finetune SimpleVQA, BVQA, FAST-VQA, and DOVER on T2VQA-DB. All results are averaged after ten-fold splitting.
@@
Tab. 2 lists the performance comparison between T2VQA and other SOTA models. Results show that T2VQA performs best in SROCC, surpassing the second by 4.68% in SROCC and 4.85% in PLCC. The zero-shot models all have relatively low scores. They either only consider the text-video alignment or don’t analyze the temporal domain information within video frames. The VQA models have higher scores, indicating that video fidelity heavily affects the assessment of text-generated video quality. However, a single perspective from video fidelity cannot address the problem properly, as there are circumstances where a high-fidelity video is generated but does not match the prompt.
Fig. 6 shows scatter plots between the predicted scores and the Z-Score MOSs of BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA. We randomly sample 1,000 videos from T2VQA-DB for testing. A better model should have a fitted curve close to the diagonal and have less dispersed scatter points. As shown in Fig. 6, T2VQA also outperforms the others.
5.2.3. Cross-dataset Validation on Sora.
Sora (Brooks et al., 2024) has been considered the SOTA T2V generation model since its release. Since the datasets listed in Tab. 1 haven’t released their subjective scores, we collect the videos generated by Sora To validate the generalization of T2VQA and other models. We collect in total of 48 videos from the official website of Sora. We invite 20 annotators to score the quality of each video. We have all the reference models tested on the Sora videos. T2VQA and other reference VQA models are trained on T2VQA-DB and tested on Sora videos. We report the SROCC, PLCC, KRCC, and RMSE between the model predictions and the ground truth MOSs. The results are listed in Tab. 2.
Experimental results show that T2VQA has the best ability of generalization among all models. Noticed that there is a performance drop between the validation on T2VQA-DB and the testing on Sora. That’s because the Sora videos have the attributes of high resolution, high frame rate, and long length, which the other current T2V models are not able to generate. We will include Sora generated videos in T2VQA-DB in future work.
@@
We also conduct a qualitative analysis on three examples with good, fair, and poor quality. We use SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA to predict their quality. Fig. 7 presents the prompts, video frames, and model predictions on the three examples, and Tab. 3 lists the models’ predictions and MOSs. Results show that T2VQA has more subjective-aligned predictions. Fig. 7a shows an example with a relatively high score. The scene in the video matches the description in the prompt well, and the video frames are clear and consistent. Fig. 7b is a medium-level example. Though the video matches the prompt basically, the video frames suffer from blurriness, which is reflected in the MOS and T2VQA’s prediction. Fig. 7c shows the worst case among the three examples. The video fails to accurately present the description in the prompt. Besides, it loses consistency between video frames. Although it has a high definition in each frame separately, it still has low scores in both MOS and T2VQA’s prediction.
@@ 5.3. Ablation Studies
SROCC | PLCC | KRCC | RMSE | |
@@
To validate the effectiveness of each module in T2VQA, we conduct thorough ablation studies, including the alignment and fidelity encoder, the fusion module, and the regression module. Experimental results are listed in Tab. 4. All results are averaged after 10-fold splitting.
5.3.1. Alignment Encoder.
@@
In T2VQA, we use the Swin-T as the backbone of the fidelity encoder. To investigate the effectiveness of the transformer-based architecture, we conduct the control experiment by using the convolution-based 3D ResNet (Hara et al., 2018) as the fidelity encoder. The results show that Swin-T has a superior performance to ResNet, validating the effectiveness of the transformer-based architecture.
@@
6. Limitations and Future Work
We investigate the limitations in the proposed T2VQA-DB and T2VQA. models like Sora (Brooks et al., 2024) have achieved stunning results with 1080P, 60s high-fidelity videos. Other current T2V models are not capable of generating such high-fidelity videos. We will expand our dataset with Sora generated videos in future work. For T2VQA, since the datasets listed in Tab. 1 do not release their subjective scores, we will conduct the cross-dataset validation to validate the generalization of T2VQA over other T2V datasets in future work.
@@ Acknowledgements.
The work was supported in part by the National Natural Science Foundation of China under Grant 62301310 and U1908210, and in part by the Shanghai Pujiang Program under Grant 22PJ1406800.References
@@
- Li et al. (2024) Chunyi Li, Haoning Wu, Zicheng Zhang, Hongkun Hao, Kaiwei Zhang, Lei Bai, Xiaohong Liu, Xiongkuo Min, Weisi Lin, and Guangtao Zhai. 2024. Q-Refine: A Perceptual Quality Refiner for AI-Generated Image. arXiv preprint arXiv:2401.01117 (2024).

### v5

@@ 4.4. Quality Regression
LLMs have been proven to be competitive on quality assessment tasks (Wu et al., 2024b, 2023f; Zhang et al., 2023c; Wu et al., 2024c; Zhang et al., 2024a). Inspired by them, we also utilize an LLM as the quality regression module in T2VQA. We first design a text instruction prompt, i.e., “Please rate the quality of this video.”, to guide the LLM. The encoded instruction and the fused feature are concatenated as the input of the LLM. Following (Zhang et al., 2023e; Wu et al., 2023f), we supervise the LLM to output one among five ITU-standard (ITU-R, 2002) levels (bad, poor, fair, good, and excellent) to represent the quality of the videos, denoted as <level>. We assign them weights of in order. Since the logit at <level> in LLM is the probability distribution of all tokens, it can be used to represent how the LLM predicts the quality of the video. Therefore, a softmax for each token is calculated and multiplied by its weight. We have the final predicted score as:
@@ 5.1. Implement Details
is a hyper-parameter for balancing, and is set to in training.
@@
Along with PLCC, we include Spearman’s Rank-order Correlation Coefficient (SRCC), Kendall’s Rank-order Correlation Coefficient (KRCC), and Root Mean Square Error (RMSE) as performance criteria. Better models should have larger SRCC, KRCC, and PLCC scores, but conversely for RMSE. Before calculating PLCC, we follow (Group et al., 2000) to map the objective score to the subject score using a four-parameter logistic function.
@@ 5.2. Performance Comparison
We use CLIPSim (Wu et al., 2021), BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), UMTScore (Liu et al., 2024), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and Q-Align (Wu et al., 2023f) as the reference algorithms. Q-Align is considered the SOTA VQA method to date. CLIPSim averages the similarity values between the text and each video frame. We adopt the same operation on BLIP and ImageReward to tune IQA metrics into VQA metrics. ViCLIP and UMTScore are metrics designed for measuring text-video alignment. SimpleVQA, BVQA, FAST-VQA, DOVER, and Q-Align are models designed for general VQA tasks. We use the pre-trained weights of CLIPSim, BLIP, and ImageReward for zero-shot testing, since they are originally designed for generated content. SimpleVQA, BVQA, FAST-VQA, DOVER, and Q-Align are originally designed for UGC videos, so we finetune them on T2VQA-DB. All results are averaged after ten-fold splitting.
@@
The first four columns of number in Tab. 2 show the performance comparison between T2VQA and other SOTA models on T2VQA-DB. Results show that T2VQA performs best in SRCC, surpassing Q-Align by 4.79% in SRCC and 3.84% in PLCC. The zero-shot models all have relatively low scores. They either only consider the text-video alignment or don’t analyze the temporal domain information within video frames. The VQA models have higher scores, indicating that video fidelity heavily affects the assessment of text-generated video quality. However, a single perspective from video fidelity cannot address the problem properly, as there are circumstances where a high-fidelity video is generated but does not match the prompt.
Fig. 6 shows scatter plots between the predicted scores and the Z-Score MOSs of BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA. The figure uses 1,000 videos, which are randomly sampled from the testing set in one split of T2VQA-DB. A better model should have a fitted curve close to the diagonal and have less dispersed scatter points. As shown in Fig. 6, T2VQA also outperforms the others.
5.2.3. Cross-dataset Validation.
ECTV (Liu et al., 2023) and TVGE (Wu et al., 2024a) are the only two T2V datasets that have released their subjective scores. We randomly sample 800 videos from each dataset for cross-dataset validation. Besides, Sora (Brooks et al., 2024) has been considered the SOTA T2V generation model. We collect the videos generated by Sora from its official website to validate the generalization of T2VQA and other models on high-fidelity videos. We invite 20 annotators to score the quality of each Sora video. T2VQA and other reference VQA models are trained on T2VQA-DB and tested on the sampled ECTV, TVGE, and Sora videos. We report the SRCC and PLCC between the models predictions and the ground truth MOSs. The results are listed in the six columns on the right side of Tab. 2.
Experimental results show that T2VQA has the best generalization ability among all models. Noticed that there is a performance drop between the validation on T2VQA-DB and the testing on other datasets. That’s because videos from the other datasets have the attributes of high resolution, high frame rate, and long length, which current open-sourced T2V models are not able to generate. We will include videos with high fidelity in T2VQA-DB in future work.
@@
We also conduct a qualitative analysis on three examples with good, fair, and poor quality. We use SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA to predict their quality. Fig. 7 presents the prompts, video frames, and model predictions on the three examples, and Tab. 3 lists the models’ predictions and MOSs. Results show that T2VQA has more subjective-aligned predictions. Fig. 7a shows an example with a relatively high score. The scene in the video matches the description in the prompt well, and the video frames are clear and consistent. Fig. 7b is a medium-level example. Though the video matches the prompt basically, the video frames suffer from blurriness, which is reflected in the MOS and T2VQA’s prediction. Fig. 7c shows the worst case. The video fails to accurately present the description in the prompt. Besides, it loses consistency between video frames. Although it has a high definition in each frame separately, it still has low scores in both MOS and T2VQA’s prediction.
@@ 5.3. Ablation Studies
To validate the effectiveness of each module in T2VQA, we conduct thorough ablation studies, including the alignment and fidelity encoder, the fusion module, and the regression module. Results are listed in Tab. 4. All results are averaged after 10-fold splitting.
5.3.1. Alignment Encoder.
@@
| SRCC | PLCC | KRCC | RMSE | |
@@
In T2VQA, we use the Swin-T as the backbone of the fidelity encoder. To investigate the effectiveness of the transformer-based architecture, we conduct the control experiment by using the convolution-based 3D ResNet (Hara et al., 2018) as the fidelity encoder. The results show that Swin-T has a superior performance to ResNet, validating its effectiveness.
@@ Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62301310, 62225112, and U1908210, and in part by the Shanghai Pujiang Program under Grant 22PJ1406800, and in part by Sichuan Science and Technology Program under Grant 2024NSFSC1426.References
@@
- Hosu et al. (2017) Vlad Hosu, Franz Hahn, Mohsen Jenadeleh, Hanhe Lin, Hui Men, Tamás Szirányi, Shujun Li, and Dietmar Saupe. 2017. The Konstanz natural video database (KoNViD-1k). In 2017 Ninth International Conference on Quality of Multimedia Experience (QoMEX). IEEE, 1–6.
@@
- Li et al. (2024a) Chunyi Li, Tengchuan Kou, Yixuan Gao, Yuqin Cao, Wei Sun, Zicheng Zhang, Yingjie Zhou, Zhichao Zhang, Weixia Zhang, Haoning Wu, et al. 2024a. Aigiqa-20k: A large database for ai-generated image quality assessment. arXiv preprint arXiv:2404.03407 2, 3 (2024), 5.
- Li et al. (2024b) Chunyi Li, Haoning Wu, Zicheng Zhang, Hongkun Hao, Kaiwei Zhang, Lei Bai, Xiaohong Liu, Xiongkuo Min, Weisi Lin, and Guangtao Zhai. 2024b. Q-Refine: A Perceptual Quality Refiner for AI-Generated Image. arXiv preprint arXiv:2401.01117 (2024).
@@
- Yang et al. (2024) Liu Yang, Huiyu Duan, Long Teng, Yucheng Zhu, Xiaohong Liu, Menghan Hu, Xiongkuo Min, Guangtao Zhai, and Patrick Le Callet. 2024. Aigcoiqa2024: Perceptual quality assessment of ai generated omnidirectional images. arXiv preprint arXiv:2404.01024 (2024).
@@
- Zhang et al. (2024a) Zicheng Zhang, Haoning Wu, Yingjie Zhou, Chunyi Li, Wei Sun, Chaofeng Chen, Xiongkuo Min, Xiaohong Liu, Weisi Lin, and Guangtao Zhai. 2024a. LMM-PCQA: Assisting Point Cloud Quality Assessment with LMM. arXiv preprint arXiv:2404.18203 (2024).
- Zhang et al. (2024b) Zicheng Zhang, Yingjie Zhou, Chunyi Li, Kang Fu, Wei Sun, Xiaohong Liu, Xiongkuo Min, and Guangtao Zhai. 2024b. A Reduced-Reference Quality Assessment Metric for Textured Mesh Digital Humans. In ICASSP 2024-2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 2965–2969.
@@
- Zhou et al. (2024a) Xunchu Zhou, Xiaohong Liu, Yunlong Dong, Tengchuan Kou, Yixuan Gao, Zicheng Zhang, Chunyi Li, Haoning Wu, and Guangtao Zhai. 2024a. Light-VQA+: A Video Quality Assessment Model for Exposure Correction with Vision-Language Guidance. arXiv preprint arXiv:2405.03333 (2024).
- Zhou et al. (2024b) Yingjie Zhou, Zicheng Zhang, Wei Sun, Xiaohong Liu, Xiongkuo Min, Zhihua Wang, Xiao-Ping Zhang, and Guangtao Zhai. 2024b. THQA: A Perceptual Quality Assessment Database for Talking Heads. arXiv preprint arXiv:2404.09003 (2024).